Consider a simple paging system with the following parameters: 232 bytes of physical memory; page size...

  1. Consider a simple paging system with the following parameters: 232 bytes of physical memory; page size of 210 bytes; 216 pages of logical address space.
    1. How many bits are in a logical address?
    2. How many bytes in a frame?
    3. How many bits in the physical address specify the frame?
    4. How many entries in the page table?
Answer #1

given the various parameter

232 bytes of physical memory

page size of 210 bytes

216 pages of logical address space

a) no. of bits are in logical address = 10+16=26 bits

b) Bytes in a frame= 2^10 byte

c) no. of bits  in the physical address specify the frame =32-10=22

d) no. of entries in the page table = 2^16=65536 entries

:_